### v2

# Proper Name Diacritization for Arabic Wikipedia: A Benchmark Dataset

Abstract
[…]
Proper Name Diacritization for Arabic Wikipedia:
A Benchmark Dataset
Rawan Bondok1, Mayar Nassar1,2, Salam Khalifa3,1, Kurt Micallef4,1, Nizar Habash1 Computational Approaches to Modeling Language (CAMeL) Lab 1New York University Abu Dhabi, 2Ain Shams University 3Stony Brook University, 4Department of Artificial Intelligence, University of Malta {rawan.bondok,nizar.habash}@nyu.edu, mayar.nassar@art.asu.edu.eg salam.khalifa@stonybrook.edu, kurt.micallef@um.edu.mt
[…]
Proper Name Diacritization for Arabic Wikipedia:
A Benchmark Dataset
Rawan Bondok1, Mayar Nassar1,2, Salam Khalifa3,1, Kurt Micallef4,1, Nizar Habash1 Computational Approaches to Modeling Language (CAMeL) Lab 1New York University Abu Dhabi, 2Ain Shams University 3Stony Brook University, 4Department of Artificial Intelligence, University of Malta {rawan.bondok,nizar.habash}@nyu.edu, mayar.nassar@art.asu.edu.eg salam.khalifa@stonybrook.edu, kurt.micallef@um.edu.mt
1 Introduction
[…]
The work presented in this paper lies at the intersection of three commonly but often independently studied Arabic NLP tasks: transliteration, diacritization, and lemmatization. Transliteration focuses on mapping words, primarily proper names, from one script to another, usually in the context of machine translation Beesley (1997); Benites et al. (2020); Chen et al. (2018). It poses challenges due to misalignments between scripts, differences in representing phonological and morphological features, and other historical ad hoc conventions. Diacritization, or diacritic restoration, aims to recover omitted diacritical marks in languages that rely on them for disambiguation Alqahtani et al. (2019); Darwish et al. (2017); Abandah et al. (2015). While both transliteration and diacritization have been well-studied for Arabic, they are typically treated in isolation. One exception is the work of Mubarak et al. (2009), which considers both in the context of Arabic-to-English name transliteration. Lemmatization focuses on mapping inflected words to their base forms. This is particularly important for morphologically rich languages such as Arabic Roth et al. (2008). In the context of Wikipedia entries, providing the lemmas is useful to readers as it gives them grounding on how to interpret and later inflect the word forms properly. More concretely, we focus here on mapping pairs of undiacritized Arabic proper nouns and their English glosses to fully diacritized and lemmatized Arabic forms. The task can be viewed as partial transliteration, where Roman-script vowels help infer (or transliterate into) Arabic diacritical marks. For example, \<نخجوان> nxjwAn ‘Nakhchivan’ (from Figure 1) should ideally be mapped to \<نَخْجِوَان> nax.jiwaAn, rather than incorrect alternatives like \<نِخْجَوَان> nix.jawaAn, or \<نُخْجِوَان> nux.jiwaAn. We present a new dataset of 3,000 unique Arabic Wikipedia proper names annotated with gold lemma-level diacritizations. Each entry is paired with its English Wikipedia equivalent, enabling the study of joint diacritization and transliteration. We benchmark GPT-4o OpenAI et al. (2024), which shows promising results but struggles with spelling variants and ambiguity. The dataset covers a range of named entities (people, places, and organizations) and includes 3,362 total pairs to reflect multiple valid diacritizations based on gloss. Our contributions are:
The remainder of this paper is structured as follows. Section 2 outlines key Arabic linguistic aspects. Section 3 reviews related work. Sections 4 and 5 describe our dataset and annotation process. Section 6 presents evaluation results and error analysis.
2 Linguistic Background
Diacritic Example
Fatha \<بَ> ba /ba/
Damma \<بُ> bu /bu/
Kasra \<بِ> bi /bi/
Shadda \<بّ> $\sim$ /bb/
Sukun \<بْ> b. /b/
Dagger Alif \<بٰ> /ba:/
[…]
Long vowel /a/ \<بَا> baA /ba:/
Long vowel /u/ \<بُو> buw /bu:/
Long vowel /i/ \<بِي> biy /bi:/
Shadda + Long vowel /a/ \<بَّا> bbaA /bba:/
Shadda + Long vowel /u/ \<بُّو> bbuw /bbu:/
Shadda + Long vowel /i/ \<بِّي> bbiy /bbi:/
Glide w \<بَوْ> baw. /baw/
Glide y \<بَيْ> bay. /bay/
Table 1: Examples of Arabic diacritics, their transliterations, and phonological values. We exclude nunation diacritics as they are not used in our lemmas.
Input Arabic Gloss Lemma Arabic Transformation
\<الست> Alst Al-Sit \<سِتّ> $\sim$ DET → $\phi$
\<الواس> AlwAs Elvas \<إِلْوَاس> Ǎil.waAws Bare Alif → Alif Hamza
[…]
\<الغظاة> $\hbar$ Al-Ghadhah \<غَظَاة> $\gamma$$\hbar$ DET → $\phi$
\<فنزويليون> fnzwylywn Venezuelans \<فِنِزْوِيلِيّ> finiz.$\sim$ 3MP → $\phi$
\<الجيبوتيون> Aljybwtywn Djiboutians \<جِيبُوتِيّ> $\sim$ DET+3MP → $\phi$
Table 2: Examples of lemmatization transformations from Arabic input (inflected) words to canonical lemmas, with English glosses and corresponding changes.
Pronunciation Arabic Transliteration
(a) /bla:stik/ \<بْلَاسْتِك> b.laAs.tik
(b) /bila:stik/ \<بِلَاسْتِك> bilaAstik
(c) /bla:stik/ \<بْلَاسْتِيك> b.laAs.tiyk
(d) /bila:stik/ \<بِلَاسْتِيك> bilaAstiyk
(e) /bla:sti:k/ \<بْلَاسْتِيك> b.laAs.tiyk
(f) /bila:sti:k/ \<بِلَاسْتِيك> bilaAstiyk
(g) /bala:sti:k/ \<بَلَاسْتِيك> balaAstiyk
(h) /ibla:stik/ \<اِبْلَاسْتِيك> Aib.laAs.tiyk
Table 3: Variants of the pronunciation and transliteration of the Arabic word for ‘plastic’. Three basic spellings: (a-b) \<بلاستك> blAstk, (c-g) \<بلاستيك> blAstyk, and (h) \<ابلاستك> AblAstk, with various diacritizations.
2.1 Arabic Diacritization
[…]
2.2 Arabic Lemmatization
In Arabic morphology, a lemma is the canonical form (also known as citation form or base word) of a word that abstracts over its inflected variants, including gender, number, person, and case, as well as attached clitics Roth et al. (2008); Habash (2010). Table 2 shows examples of input forms and their corresponding lemmas. In our context, lemmatization is simpler than in free-form text: we focus only on proper nouns, an English gloss is available to guide vowelization, and clitics are rare aside from the definite article (DET) \<ال> Al. The main challenges are distinguishing between baseword and determiner uses of \<ال> Al initial substring (see Table 2 rows 1-2), and handling plural endings (3MP) in demonyms (Table 2 rows 5-6).
2.3 Arabic Transliterations
Transliteration from Roman script to Arabic presents several challenges, primarily due to the misalignment between the phonological values of the original language and its Roman script orthography, as well as differences between the phonologies of the original languages and Arabic. Arabic, for example, has fewer vowel sounds (6 in Arabic vs. 15 in English), and some missing (no /p/ or /v/) and additional consonants (e.g., emphatic /d/ and /q/). Arabic dialects vary in phonology, including sound quality, letter mapping, and syllabification, leading to multiple valid transliterations. For instance, the borrowed word ‘plastic’ can have different pronunciations and spellings, reflecting variations in vowels and syllabification (see Table 3). During annotation, we adhered to the Wikipedia spelling and aligned closely with the English gloss. The team included Egyptian and Levantine speakers, with the primary annotator being Egyptian.
3 Related Work
3.1 Diacritization in Arabic NLP
Arabic diacritization has been extensively studied using both statistical and neural methods. Some approaches treat it as a standalone task (Zitouni et al., 2006; Mubarak et al., 2019), while others integrate it into multi-task learning frameworks alongside linguistically related tasks such as part-of-speech tagging (Habash and Rambow, 2005; Alqahtani et al., 2020). A commonly adopted strategy involves the use of morphological analyzers. For instance, Camelira implements an analyze-and-disambiguate pipeline: a morphological analyzer generates candidate analyses, which are then ranked by a classifier Obeid et al. (2022). Similarly, Farasa uses morphological patterns to generate the diacritized words Darwish et al. (2017). Systems like Farasa (Darwish et al., 2017) and Camelira Obeid et al. (2022) have demonstrated strong performance sentence-level diacritization tasks. These systems, however, are not directly applicable to our task, which centers on isolated proper nouns, adheres to a task-specific diacritization schema, and incorporates lemma mapping. Unlike sentence-based systems that leverage surrounding context for disambiguation, our task involves context-free diacritization, which poses distinct challenges (see Section 5.1).
3.2 Lemmatization in Arabic NLP
Lemmatization is another core task in Arabic NLP, and several tools offer robust performance across a variety of syntactic categories (Jarrar et al., 2024; Obeid et al., 2020, 2022). However, our lemmatization task has a narrower scope: it is limited to proper nouns that have a limited inflectional space (see Section 5.1 for further details on our lemmatization space).
3.3 Transliteration in Arabic NLP
Earlier research on Arabic–English transliteration relied on statistical approaches (Abduljaleel and Larkey, 2004), followed by more targeted work on proper nouns using models such as phonemic memory networks Tian et al. (2022). A persistent challenge in this area is the lack of standardization in transliterating foreign names into Arabic, a problem exacerbated by the omission of diacritics Aziz (1983); Odisho (1992). To address the lack of standardization and limited resources, we introduce a new dataset and annotation guidelines specifically designed for the task of utilizing proper noun transliteration as a signal for Arabic Diacritization. Prior efforts investigated the intersection of transliteration and diacritization, such as Mubarak et al. (2009) and Darwish et al. (2017). Mubarak et al. (2009) used diacritization as a preprocessing step to transliteration. Although, the approach presented in Darwish et al. (2017) for automatically diacritizing transliterated words included leveraging English transliterations to generate Arabic diacritized proper nouns, both their training and test sets were limited in size (500 and 200 instances, respectively). Our resource, by contrast, is publicly available, much larger (3,000 diacritized lemmas), and benchmarked for robust evaluation and development.
3.4 Arabic Proper Noun Resources
Although various Arabic proper noun datasets exist, they often suffer from limited accessibility, lack of diacritics, or domain constraints. For example, Matthews (2007) compiled a list of 10,001 Arabic names, but the dataset is not publicly available. Eryani and Habash (2021) provide automatically Romanized Arabic bibliographic entries without diacritics, and both the Dan database (Halpern et al., 2009) and SAMA Graff et al. (2009) include diacritized proper nouns, but they were mainly collected form news sources. Khairallah et al. (2024) released a large set of proper nouns as part of their CamelMorph Arabic morphological analyzer (henceforth AraProp). The dataset consists of two distinct portions: (a) AraPropComm, which extends the SAMA Graff et al. (2009) proper-noun list and updates their diacritizations; and (b) AraPropWiki which comprises 63K entries extracted from a Wikidata dump (14-Mar-2023). [^3] https://dumps.wikimedia.org/wikidatawiki/entities/ The AraPropWiki was filtered by Khairallah et al. (2024) to include only single word entities in Arabic and English, and covering only personal and family names, locations and organizations. Unfortunately, Khairallah et al. (2024) did not provide diacritizations for the AraPropWiki portion. Our interest in this topic started by this problem in their open-source resource, which was not usable for our purposes. We discuss these datasets further in Section 4. In this work, we present the first publicly available dataset of maximally manually diacritized and lemmatized Arabic proper nouns on a portion of the AraPropWiki dataset sourced from Wikimedia and manually annotated using English equivalents in a consistent and standardized annotation scheme. To support future work, we also release detailed annotation guidelines and provide the first benchmark of GPT-4’s performance on this task, offering a new resource for evaluating Arabic proper noun diacritization and transliteration.
AraPropComm AraPropWiki AraPropWikiD3K
Unique Arabic 6,022 63,417 3,000
Arabic-English Entries 7,202 71,251 3,362
English glosses per entry 1.20 1.12 1.12
Diacritizations Yes No Yes
Table 4: Comparison of dataset statistics across AraPropComm, AraPropWiki, and the annotated subset AraPropWikiD3K.
AraPropComm AraPropWikiD3K
Entries 7,202 3,362
Avg Freq 205,077 61,544
Median Freq 11,732 75
Avg Freeman Similarity Score $\uparrow$ 0.92 $\uparrow$ 0.91
Table 5: Statistics comparing AraPropComm and AraPropWikiD3K subsets. In terms of average frequency (Avg Freq) and median frequency (Med Freq) measured against a frequency list, and Freeman Similarity Score.
4 Datasets
We work with the AraProp dataset, released as part of an Arabic morphological analyzer by Khairallah et al. (2024). As noted in Section 3, it consists of two parts: AraPropComm and AraPropWiki. We randomly selected 3,000 unique Arabic-script proper nouns from AraPropWiki for manual annotation, forming our dataset AraPropWikiD3K. Table 4 compares the three datasets in terms of unique Arabic entries and full Arabic–English gloss pairs. The original data included multiple glosses per Arabic word (12–20% extra on average). We normalized this by splitting them into separate one-to-one pairs. For example, \<آنا> ĀnA, glossed as ‘A’ana; Ana; Anna’, became three distinct entries: (\<آنا> ĀnA, ‘A’ana’), (\<آنا> ĀnA, ‘Ana’), and (\<آنا> ĀnA, ‘Anna’). Thus, our 3,000 Arabic words expanded to 3,362 Arabic–gloss pairs. Table 5 presents some additional details comparing AraPropComm and AraPropWikiD3K in terms of average and median frequency and Arabic-English phonological similarity. For frequency we used the Arabic Frequency list from Khalifa et al. (2021). For phonological similarity, we used the Freeman similarity score Freeman et al. (2006). While phonological similarity is only slightly lower in AraPropWikiD3K, the overall frequency is significantly lower, highlighting the importance of modeling the diacritization of low-frequency proper nouns in Wikipedia and NLP.
Invalid Lemma Gloss Issue Corrected Lemma
\<سَانْشِيز> sAnšiyz Sanchez Long vowels require preceding diacritics \<سَانْشِيز> saAn.šiyz
\<كَرَمُ> karamu Karam Final letter cannot have a diacritic \<كَرَم> karam
[…]
Table 6: Examples of malformed words and their corrected lemmas with transliterations.
5 Data Annotation
In this section, we discuss the diacritization guidelines we used, as well the setup for initial automatic processing followed by manual correction.
5.1 Diacritization Guidelines
We follow the Arabic maximal diacritization guidelines as presented in Elgamal et al. (2024) with a small number of modifications to fit the purpose of our task. We list the most important decisions that are different from standard Arabic diacritization.
The Lemmatization Requirement
This effort focuses exclusively on the diacritization of proper nouns and mapping them to their lemmas. As such we require the removal of clitics such as the definite article and the removal of plural suffixes (see Section 2.2).
Input Spelling Integrity
Aside from the minimal changes connected to lemmatization, and corrections of the obligatory Hamza diacritic in Alif Hamza forms (see Section 2.1), we do not add, remove, or modify any letters in the provided input.
Consonant Clusters in Foreign Names
While standard Arabic generally avoids consonant clusters, our dataset includes many foreign proper nouns where such clusters are phonetically natural. To more faithfully capture their pronunciation, we allowed forms with consecutive consonants, either multiple letters marked with Sukuns, or a Sukun followed by a letter with Shadda (geminated), even though this departs from Standard Arabic diacritization norms. For example, \<إلكتريك> Ǎlktryk ‘Electric’ should be diacritized as \<إِلِكْتْرِيك> Ǎilik.t.riyk (with the consonant cluster /tr/), and \<زدينيك> zdynyk ‘Zdeněk’ should be diacrtized as \<زْدِنِيك> z.diniyk with initial /zd/ cluster.
Final Letter Ya
The final letter \<ي> y has multiple diacritizations that intersect with changes in dialectal Arabic final y-gemination softening (into a /i/). As such, we had to dedicate part of the guidelines to outline the rules for diacritizing it as a geminated /yy/, a long vowel /i:/ or a glide /ay/. The geminated version is the most specific in requirements with three possible cases:
  • The gemination comes from the root or pattern of the word such as the final Ya in \<رَخِيّ> $\sim$ ‘Ar-Rakhi’.
  • The lemma can be interpreted as having the derivational attribution suffix Ya-Nisba, e.g., \<إِشْبِيلِيّ> Ǎiš$\sim$ ‘Sevillian’ (of or related to \<إِشْبِيلِيَّة> $\hbar$ ‘Seville’).
  • Gemination is necessary to reflect the pronunciation of certain foreign names, such as \<أَرْكْوِيّ> Âar.k.$\sim$ ‘Arcueil’.
For other cases, if the final vowel sounds like a short /i/ or a long /i:/ and has a corresponding \<ي> y, it is diacritized resembling a long vowel, e.g., \<أغاسي> ÂgAsy Agassi, should be diacritized as \<أَغَاسِي> ÂagaAsiy. The glide version is straightforward as it has a distinct phonological signal. One example is the word \<نَي> nay ‘Ney’.
Checking Well-formedness
To ensure consistency with our annotation guidelines, we implemented automated checks to validate the well-formedness of diacritized lemmas. While these checks do not guarantee correctness, they are effective at identifying common errors and inconsistencies. We use these checks on both human and automatic annotations. See Table 6 for examples.
5.2 Initial Automatic Diacritization
To speed up the annotation process, we gave our annotator an automatically diacirtized version of the data. We used GPT-4o with Arabic Input and English Gloss (comparable to the best setting in Section 6). At the time of creating the initial automatic diacritization, we assumed this a reasonable starting point.
GPT-4o postprocessing
The output of GPT-4o was not always usable as is. When applying well-formedness checks to the diacritized outputs generated by GPT-4o, we observed several recurring patterns of errors that compromised the validity of the diacritized forms. In response, we developed an automated pipeline specifically aimed at correcting these systematic errors. The automatic correction procedures included the following operations:
  • Insertion of Fatha before Alif (\<ا> A).
  • Insertion of Kasra after Alif-Hamza-Below (\<إ> Ǎ).
  • Normalization of Shadda-Vowel clusters such that the vowel diacritic follows the Shadda diacritic.
  • Removal of final diacritics as lemmas do not have them.
  • Insertion of missing Sukuns to indicate vowel absence at the end of syllable or in a consoant cluster.
  • Removal of Fatha after Alif Madda (\<آ> Ā).
  • Mapping Non-Arabic Arabic-script letters, such as those used in Urdu or Persian, to their closest Arabic language form.
5.3 Manual Diacritization
The manual diacritization and quality checks were carried out by a native speaker of Arabic who is a trained linguist and highly experienced annotator. The annotation process initially was done in tandem with the finalization of the guideline with a team of the authors working jointly to optimize the quality of the annotation. The annotator was provided an Arabic word, along with its English gloss, and a proposed diacritization from GPT-4o after being refined by the automatic post-process described above. The annotations were carried on Google Sheets in a very simple setup. The annotator reviewed the proposed diacritization making changes where needed in accordance to the guidelines. The annotator made changes to 909 proposed lemmas out of 3,362 ($\sim$27%). In 213 instances (6.3% of all entries), there was a change connected with lemmatization: 74% relative involved the Al determiner, 22.5% a change in Alif-Hamza spelling, and 3.3% involving the demonym plural ending.
6 Evaluation
6.1 Experimental Setup
We perform computational experiments to perform the task of diacritization of proper nouns. For this, we prompt GPT-4o on all of the annotated dataset described in Section 5. We prompt the model with different input formats to assess its capabilities while giving it different levels of information: the inputs and the number of examples shown to the model (shots).
Inputs
The model is given a detailed description of the task to be performed. Our main experiments reflect all the information given to our annotator, where we provide the model with both the Arabic Input and the English Gloss (Arabic Input and Gloss). Additionally, we also experiment with a more constrained setup where the model is provided solely with the the Arabic Input (Arabic Input Only).
Shots
In addition to the different inputs, we also consider further experiments where we supply the model with varying number of examples to learn from. Hence, in addition to just providing the input to diacritize (Zero-Shot), we also supply the model with a single example (One-Shot), and 80 examples (Few-Shot). The examples are randomly sampled from the AraPropComm data. However, since AraPropComm has fully lemmatized Arabic Inputs, we manually manipulated some of the examples to have a representation of clitic removal and Hamza normalization. Refer to Appendix A for a more detailed description of the prompts used. As a post-processing step, the outputs were ran through the same processing pipeline mentioned in Section 5.2. To evaluate the performance of the different experiments, we compute accuracy by measuring the exact match between the post-processed output and the gold-standard diacritization.
6.2 Results
Table 7 shows the results with different prompts.
Input Format Shots Accuracy
Arabic Input and Gloss Zero 46.49%
Arabic Input and Gloss One 61.90%
Arabic Input and Gloss Few 72.96%
Arabic Input Only Zero 36.67%
Arabic Input Only One 49.73%
Arabic Input Only Few 55.86%
Table 7: GPT-4o model Results measured by exact match accuracy.
Input Gloss Reference Prediction Error Type
[…]
\<أفراموفو> ÂfrAmwfw Avramovo \<أَفْرَامُوفُو> Âaf.raAmuwfuw \<أَفْرَامُوفُو> Âaf.raAmuwfuw Exact Match
[…]
\<إشتهارد> ĂšthArd Eshtehard \<إِشْتِهَارْد> Ăiš.tihaAr.d \<إِشْتَهَارْد> Ăiš.tahaAr.d Diac
\<بلاجيفيتش> blAjyfytš Blažević \<بْلَاجِيفِيتْش> b.laAjiyfiyt.š \<بِلَاجِيفِيتْش> bilaAjiyfiyt.š Diac
\<دسوق> dswq Desouk \<دِسُوق> disuwq \<دُسُوق> dusuwq Diac
\<ريبلاي> ryblAy Ripley \<رِيبْلَاي> riyb.laAy \<رِيبْلِي> riyb.liy AWY
[…]
\<جوندريزيك> jwndryzyk Gondrezick \<جُونْدْرِيزِيك> juwn.d.riyziyk \<جُنْدْرِيزِيك> jun.d.riyziyk AWY
[…]
\<تسيخانوف> tsyxAnwf Ciechanów \<تِسِيخَانُوف> tisiyxaAnuwf \<تِسِيهَانُوف> tisiyhaAnuwf h $\leftrightarrow$ x
\<أردينة> $\hbar$ Ardineh \<أَرْدِينَة> $\hbar$<أَرْدِينَه> Âar.diynah $\hbar$leftrightarrow h
[…]
\<كرامة> $\hbar$ Gourrama \<كُرَامَة> $\hbar$<كُورَامَا> kuwraAmaA Multiple
\<ايكوميديا> AykwmydyA Eco-Médias \<إِيكُومِيدْيَا> Ăiykuwmiyd.yaA \<إِيكُومِيدْيَآس> Ăiykuwmiyd.yaĀs Multiple
\<بارافرانكا> bArAfrAnkA Barrafranca \<بَارَّافْرَانْكَا> baAr aAf.raAn.kaA \<بَارَافْرَانْكَة> $\hbar$ Multiple
Table 8: Examples of evaluated instances along with their, reference and predicted diacritized forms, and corresponding error types. The error categories are diacritic mismatches (Diac), AWY spelling changes (AWY), several consonant and ta-marbuta substitutions (j $\leftrightarrow$ $\gamma$, h $\leftrightarrow$ x, and $\hbar$leftrightarrow h), and those with multiple changes (Multiple).
The results demonstrate that while diacritizing proper nouns remains a challenging task, incorporating the English gloss offers a valuable signal for the model. Notably, the best performance is achieved with few-shot prompting, showing the effectiveness of providing a diverse and representative sample.
6.3 Error Analysis
We analyzed errors from a sample of 1,010 output entries from the best performing setup from Section 6.2, and classified errors into several categories based on observed patterns. There were 740 (73.3%) exact matches (correct generations). Of the 270 (26.7%) errors, there were 175 cases where the error was only diacritization differences. See examples in Table 8. Upon further analysis of this class of errors, we found that the model overpredicts fathas (+25%) and shaddas (+96%), while underpredicting kasras (-18%) and sukuns (-23%), indicating imbalanced vowel modeling and overuse of gemination. The next largest class of errors, 60 cases, where those with spelling changes limited to the set of long vowel (and glides) letters \<ا> A, \<و> w, and \<ي> y (AWY). As we see in the examples in Table 8, the model has the tendency of dropping such letters rather than adding them. Another class of errors, 10 cases, are those with specific letter replacements such as \<ج> j $\leftrightarrow$ \<غ> $\gamma$, \<خ> x $\leftrightarrow$ \<ه> h, and \<ـة> $\hbar$leftrightarrow \<ـه> h. The final class of errors, 25 cases, are those with multiple changes happening at once. While these cases don’t match the gold reference, they are plausible and acceptable alternatives in most cases, especially in the context of linguistic variation discussed in Section 2. For example, the generated diacritization for \<بلاجيفيتش> blAjyfytš ‘Blažević’ as seen in Table 8, follows the common phenomena of breaking word initial complex onsets found in many spoken dialects of Arabic and in MSA. Another example is the entry \<ايكوميديا> AykwmydyA ‘Eco-Médias’, where the input follows a pronunciation-based transliteration while the generated form adhered to the orthography of the gloss. These observations indicate the need for exploring different modeling techniques that accommodate this special case of Arabic diacritization which in turn requires additional annotated examples.
7 Conclusion and Future Work
We presented a new 3,362-entry dataset of Arabic Wikipedia proper names annotated with gold-standard lemma diacritizations, paired with their English equivalents. This resource enables the joint study of diacritization and transliteration in a realistic setting characterized by ambiguity and spelling variation. We benchmarked GPT-4o on this task, providing insights into its capabilities and limitations. While the model performs reasonably well, especially on frequent names, it struggles with rarer entries and variant mappings. Looking ahead, we plan to expand the dataset with more diverse names, integrate it into a morphological analyzer, and explore fine-tuned models for diacritizing proper nouns in broader contexts. We also plan to fine-tune dedicated models for this task and develop more robust approaches to name ambiguity, especially with multiple valid diacritizations. We hope this resource advances Arabic NLP and name normalization in multilingual settings like Wikipedia.
Limitations
A primary limitation of this work lies in the inherent subjectivity of diacritization, particularly for proper nouns where multiple correct variants may exist depending on regional, historical, or phonetic conventions. Despite rigorous annotation guidelines and quality checks, variability is an inevitable aspect of any human-annotated linguistic resource. Our current benchmark relies solely on GPT-4o, and we acknowledge the importance of evaluating performance across a broader range of large language models. While initial results are promising, the overall performance remains limited and, in our assessment, not yet suitable for reliable downstream use.
Ethics Statement
All data used in this project were sourced from publicly available Arabic Wikipedia entries and their corresponding English titles, in accordance with Wikimedia’s terms of use. The annotation process was conducted transparently and ethically, with fair compensation provided to the annotator. We make both the corpus and the annotation guidelines publicly accessible under an open license, supporting reproducibility and community collaboration. Our goal is to contribute a valuable resource for Arabic language processing and to aid the broader Wikimedia effort by enhancing the quality of Arabic Wikipedia entries. Finally, we acknowledge that all NLP tools and resources can be used with malicious intent; this is not our intention, and we categorically discourage it.
Benefits
This work directly supports the Wikimedia community by enhancing the quality and accessibility of Arabic Wikipedia content. By providing more accurate diacritization for proper nouns from all over the world on Arabic Wikipedia, we aim to improve readability, pronunciation, and downstream tasks such as named entity recognition and machine translation. The dataset, code, and annotation guidelines are all released under the Creative Commons Attribution-ShareAlike (CC BY-SA) license to ensure community reuse and adaptation. Filtering was applied to select single-word proper nouns related to people, locations, and organizations, drawn from Arabic Wikipedia entries that have clear English counterparts, thereby supporting multilingual alignment and cross-lingual research.
Risks
Our project poses no known risks to Wikimedia editors or contributors. We do not name, identify, or reference any individual editor (by username or otherwise), nor do we expose any metadata that could be used to infer editor identities. The work focuses solely on content-level linguistic annotation and transformation. There are no known ways in which this research could be used to derive sensitive or personal information about contributors, and we strongly discourage any attempts to repurpose the resource for such purposes.
Appendix A GPT-4o Prompts
In the system role, we provide the task description, and optionally, the few-shot demonstrations, when they are used. For the user role, we always provide the single instance to be diacritized. Table 9 lists all of the prompts used for the different settings.
Shots Prompt
Arabic Word+Gloss Input
Zero You are an expert in Arabic. You are given the undiacritized proper noun in Arabic and its English gloss. Your task is to generate the corresponding diacritized proper noun lemma in Arabic. Arabic lemmas are dictionary entries that have no attached definite article (\<ال>). Diacritization is adding the correct diacritic markings to undiacritized words. Remove the Arabic definite article (\<ال>) when present. Do not add, remove, or substitute any other letters in the input. Determine the most accurate diacritization that matches the English gloss pronunciation. The user will provide a Markdown table with 1 rows. Each row contains an undiacritized proper noun in Arabic in the “Input” column and its English gloss in the “Gloss” column. Return exactly 1 diacritized lemmas, one per line. Do not include extra text, explanations, or formatting.
Few/One You are an expert in Arabic. You are given the undiacritized proper noun in Arabic and its English gloss. Your task is to generate the corresponding diacritized proper noun lemma in Arabic. Arabic lemmas are dictionary entries that have no attached definite article (\<ال>). Diacritization is adding the correct diacritic markings to undiacritized words. Remove the Arabic definite article (\<ال>) when present. Do not add, remove, or substitute any other letters in the input. Determine the most accurate diacritization that matches the English gloss pronunciation. The user will provide a Markdown table with 1 rows. Each row contains an undiacritized proper noun in Arabic in the “Input” column and its English gloss in the “Gloss” column. Return exactly 1 diacritized lemmas, one per line. Do not include extra text, explanations, or formatting. Here are some examples of triplets of an undiacritized proper noun in Arabic (“Input”), its respective English gloss (“Gloss”), and its diacritized lemma (“Output”) for reference <Few-Shots-table>
Arabic Word Only Input
Zero You are an expert in Arabic. You are given the undiacritized proper noun in Arabic. Your task is to generate the corresponding diacritized proper noun lemma in Arabic. Arabic lemmas are dictionary entries that have no attached definite article (\<ال>). Diacritization is adding the correct diacritic markings to undiacritized words. Remove the Arabic definite article (\<ال>) when present. Do not add, remove, or substitute any other letters in the input. The user will provide a Markdown table with 1 rows. Each row contains an undiacritized proper noun in Arabic in the “Input” column. Return exactly 1 diacritized lemmas, one per line. Do not include extra text, explanations, or formatting.
Few/One You are an expert in Arabic. You are given the undiacritized proper noun in Arabic. Your task is to generate the corresponding diacritized proper noun lemma in Arabic. Arabic lemmas are dictionary entries that have no attached definite article (\<ال>). Diacritization is adding the correct diacritic markings to undiacritized words. Remove the Arabic definite article (\<ال>) when present. Do not add, remove, or substitute any other letters in the input. The user will provide a Markdown table with 1 rows. Each row contains an undiacritized proper noun in Arabic in the “Input” column. Return exactly 1 diacritized lemmas, one per line. Do not include extra text, explanations, or formatting. Here are some examples of pairs of an undiacritized proper noun in Arabic (“Input”), and its diacritized lemma (“Output”) for reference <Few-Shots-table>
Table 9: System prompts used in the experiments. <Few-Shots-table> is a placeholder for few-shot examples. In either setting, the user prompts consist solely of a single instance to be diacritized.
Table 10 shows a sample of the few-shot examples. These are formatted as a markdown table in the prompts.
Arabic Word Gloss Diacritized Reference
\<الاستانة> $\hbar$ Istana \<إِسْتَانَة> $\hbar$
\<ايدكس> Aydks IDEX \<إِيدِكس> Ǎiydiks
\<الغارديان> AlgArdyAn Guardian \<غَارْدِيَان> gaAr.diyaAn
\<رودريغيز> rwdrygyz Rodriguez \<رُودْرِيغِيز> ruwd.riygiyz
\<اوروغواي> AwrwgwAy Uruguay \<أُورُوغْوَاي> Âuwruwg.waAy
\<بوتيه> bwtyh Boutier \<بُوتِيِه> buwtiyih
\<وايزمن> wAyzmn Weizman \<وَايزْمَن> waAyz.man
Table 10: A sample of few-shot examples used for prompting GPT-4o